Determine the adequacy of Bolt B when 20mm diameter A325 bolts are used with threaded excluded from the shear planes. Assume the strength of the column and the ST section does not govern. The value of F is 500kN, with a 60-degree angle from the horizontal. Bolts A F Bolts B


1
Expert's answer
2021-07-22T14:11:01-0400

For ASTM A325= yield strength "= 92 ksi =634.318 N\/mm^2"

Shear stress acting on bolts "F= 500 cos 60=250 kN"

"A=2 (\frac{\pi}{4}*20^2) =628.3 mm^2"

Shear stress due to load= "\frac{250000}{628.3}= 397.89 N/mm2"

Nominal strength for two bolts > shear stress acting on the bolts

"2* 634.318 N\/mm ^2>397.89 N\/mm^2"

"1268.636 N\/mm^2 >397.89 N\/mm^2"

Therefore bolts are adequate